# Multiscale Quantile Regression with Local Error Control

### v2

@@ Appendix C Implementation details
4:$j_{\min}\leftarrow 1$, $J\leftarrow 0$, $C_{1,n}=\infty$, $\widehat{f}_{1,1}\leftarrow Z_{1}$, $C_{1,1}=0$ and $J_{1}\leftarrow 0$
@@
15:                     for $l=j-i,\dots,1$ do $\triangleright$ If using dyadic intervals, $l=2^{p},2^{p-1},\dots,1$ with $p=\lfloor{{\log_{2}(j-i)}}\rfloor$
@@
32:                             $S\leftarrow 1$triangleright Pruning 1: as no $j^{\prime}>j$ with $\mathcal{I}([i/n,j^{\prime}/n])=1$, all indices $j^{\prime}>j$ can be dropped
@@
35:                             break $\triangleright$ Pruning 2: as no $i^{\prime}<i$ with $\mathcal{I}([i^{\prime}/n,j/n])=1$, all indices $i^{\prime}<i$ can be dropped
@@ C.1 Computation of local tests
$T_{I}(Z,\theta)=\underset{J\subseteq I}{\max}\sqrt{2L_{J}(Z,\theta)}-\sqrt{2\log\left(\frac{e\lvert I\rvert}{\lvert J\rvert}\right)},$
where
$L_{J}(Z,\theta)$ $=\lvert J\rvert\left[\overline{W}_{J}\log\left(\frac{\overline{W}_{J}}{\beta}\right)+(1-\overline{W}_{J})\log\left(\frac{1-\overline{W}_{J}}{1-\beta}\right)\right],$
$\overline{W}_{J}$ $=\frac{1}{\lvert J\rvert}\sum_{i\in J}W_{i},$
$W_{i}$ $=W_{i}(Z_{i},\theta)=\mathbbm{1}_{\{Z_{i}\leq\theta\}}.$
Note that the multiscale constraint $T_{I}(Z,\theta)\leq q_{\alpha}(\lvert I\rvert)$ is equivalent to for all subintervals $J\subseteq I$,
$\overline{W}_{J}\log\left(\frac{\overline{W}_{J}}{\beta}\right)+(1-\overline{W}_{J})\log\left(\frac{1-\overline{W}_{J}}{1-\beta}\right)\;\leq\;\frac{\left(q_{\alpha}(\lvert I\rvert)+\sqrt{2\log\frac{e\lvert I\rvert}{\lvert J\rvert}}\right)^{2}}{2\lvert J\rvert}\;\eqqcolon\;\tilde{q}_{\lvert I\rvert,\lvert J\rvert}.$ (26)
Here $\tilde{q}_{\lvert I\rvert,\lvert J\rvert}=\tilde{q}_{k,l}$ depends only on $k=\lvert I\rvert$ and $l=\lvert J\rvert$. The function $g_{\beta}:(0,1)\to\mathbb{R}$
$g_{\beta}:\quad x\mapsto x\log\left(\frac{x}{\beta}\right)+(1-x)\log\left(\frac{1-x}{1-\beta}\right)$
is strictly convex on $(0,1)$ and has a unique minimizer at $x=\beta$. Let $u_{k,l}$ and $v_{k,l}$ with $u_{k,l}\leq v_{k,l}$ be the solutions of $g_{\beta}(x)=\tilde{q}_{k,l}$. Then the inequality in (26) becomes
$\overline{W}_{J}\in[u_{k,l},v_{k,l}],\quad\text{for all subintervals }J\subseteq I,\,k=\lvert I\rvert\text{ and }l=\lvert J\rvert.$
By the definition of $W_{i}$’s, the above displayed relation holds if and only if
$\theta\in\bigcap_{J\subseteq I}\left[Z_{J,u_{k,l}},Z_{J,v_{k,l}}\right]$ (27)
where $Z_{J,q}$ denotes the $q$-quantile of $Z_{J}=(Z_{i})_{i\in J}$ for any $q\in(0,1)$. The intersection on the right hand-side of (27) is empty if and only if there is no $\theta\in\mathbb{R}$ fulfils the multiscale constraints on $I$, in which case at least one additional change point is required. Note that the subinterval $J$ is varying in $I$, and the corresponding quantities $u_{k,l}$ and $v_{k,l}$ are changing as well. It amounts to computing the so-called range quantiles, which is formulated as the range quantile problem below.
@@ 2. Output:
In computer science the range quantile problem is solved in two steps: a preprocessing step and a query step. The preprocessing step transforms the original input $Y$ into a certain data structure, and in the query step each query is answered efficiently based on the data structure created in the preprocessing step. We focus on the data structure of wavelet tree and refer to Remark 5 for alternative data structures.
A wavelet tree is a binary tree whose nodes consist of a value vector and a rank vector. The value vector of root is exactly the whole input and it is spitted into two children by comparing with the median of the value vector. More precisely, its left child contains all elements smaller than median while its right child consists of all elements larger or equal than median (for simplicity we assume that all $n$ data are distinct). For $i=1,\dots,n$, the $i$-th entry of rank vector in root indicates the number of elements belonging to the left child between the first and $i$-th value. A wavelet tree can be constructed recursively until its leaves (i.e., nodes with only one element). Note that it is possible to compute all splitting required medians in $O(n\log n)$ steps and each layer can be constructed in $O(n)$ time. Thus, a wavelet tree is created in $O(n\log n)$ steps. Each single query $Q=(L,R,q)$ is answered in $O(\log n)$ steps as follows: Let $k=\lceil{(R-L+1)\cdot q}\rceil$ denote order of the $q$-th quantile of $(Y_{L},\dots,Y_{R})$. At each node we compute the difference between the rank of $R$ and rank of $L-1$, which equals to the number of elements in $(Y_{L},\dots,Y_{R})$ belonging to its left child. If it is greater or equal than $k$, then the target value is in the left child; otherwise, the target value is in the right child. We continue this process until the last layer and return value in the last leaf. The computation cost of at each node is $O(1)$, since all ranks are precomputed. Note the depth of a wavelet tree is $O(\log n)$, so the computational complexity of each query is $O(\log n)$.
We summarise these results and obtain the following proposition.
Proposition 12.
C.2 Proof of Theorem 6
Now we are ready to prove Theorem 6 and proceed in the following steps.
Recursion. Consider the recursive relation in (16). The subproblem with $j=1$ is solved immediately (cf. line 1 in Algorithm 1). For any $j\geq 2$, the $j$-th subproblem can be solved recursively by (16) via checking whether there is a constant function fulfilling the multiscale on $[i/n,j/n)$, i.e., $\mathcal{I}([i/n,j/n))=1$ (cf. lines 12-27 in Algorithm 1). One should keep in mind that even if we knew $\mathcal{I}\left(\left[i/n,j/n\right)\right)=0$ for some $i$, we can not conclude $\mathcal{I}\left(\left[i^{\prime}/n,j/n\right)\right)=0$ for $i^{\prime}<i$, due to the varying penalties and quantiles. Since $\widehat{K}[j]$ requires the minimisation among all indices $i\leq j-1$ with $\mathcal{I}([i/n,j/n))=1$, all indices should be taken into account.
Pruning. Nevertheless, the checking of some indices can be avoided if we employ some pruning techniques similarly as in Li et al., (2016). Towards this, we consider the test statistic $T_{I}^{0}(Z,\theta)={\max}_{J\subseteq I}\sqrt{2L_{J}(W,\theta)}-\sqrt{2\log\left({en}/{\lvert J\rvert}\right)}$, where the largest penalty with size $n$ is subtracted. Then it holds for all $Z$, $\theta$ and $I$ that $T_{I}(Z,\theta)\geq T_{I}^{0}(Z,\theta)$. During the minimisation over indices ranging from $i=j-1$ to $i=0$, if we find an index $i$ with $\mathcal{I}\left(\left[i/n,j/n\right)\right)=0$, then we check whether there exists a $\theta\in\mathbb{R}$ such that
$T_{[i/n,j/n]}^{0}(Z,\theta)\leq\max_{m\leq n}q_{\alpha}(m),$
which is the most relaxed multiscale constraint. If no constant $\theta$ fulfils this constraint displayed above, then $\mathcal{I}(I)=0$ for all $I$ containing $[i/n,j/n)$ and thus the indices $11,\dots,i-$ 1 can be excluded from further consideration (cf. lines 29-33 in Algorithm 1).
Final dynamic program. We formalise the final procedure as follows. Let $\mathcal{A}_{0}=\{0\}$ and $\mathcal{B}_{0}=\{1,\dots,n\}$ and for $k=1\dots n$,
[…]
$\mathcal{A}_{k}$$\coloneqq\{j\in\mathcal{B}_{k-1}\cap[1,r_{k}]:\,\mathcal{I}([i/n,j/n))=1\text{ for some }i\in\mathcal{A}_{k-1}\},$
$\mathcal{B}_{k}$ $\coloneqq\mathcal{B}_{k-1}\setminus\mathcal{A}_{k}.$
Complexity analysis. In the computation of MUSCLE, we first preprocess data $(Z_{1},\dots,Z_{n})$ by constructing a wavelet tree, which requires $O(n\log n)$ steps and $O(n\log n)$ memory (cf. line 1 in Algorithm 1). Then, by Proposition 12, $\mathcal{I}([i/n,j/n))$ can be computed in $O((j-i)^{2}\log n)$ steps (cf. lines 13-18 in Algorithm 1). Thus, the overall complexity for computing $\widehat{K}$ is bounded from above by
$O\left(\sum_{k=0}^{\widehat{K}}\sum_{i\in\mathcal{A}_{k}}\sum_{j=i+1}^{r_{k+1}}(j-i)^{2}\log n\right)\leq O\left(\log n\sum_{k=0}^{\widehat{K}}\lvert\mathcal{A}_{k}\rvert\left(r_{k+1}-\min\mathcal{A}_{k}\right)^{3}\right)\\ \leq O\left(n\log n\cdot\underset{0\leq k\leq\widehat{K}}{\max}\left(r_{k+1}-\min\mathcal{A}_{k}\right)^{3}\right).$ (28)
Furthermore, the MUSCLE estimator $\widehat{f}$ in (7) is computed in the meanwhile of calculating $\widehat{K}$. More precisely, when checking each index in the dynamic program, we need only $O(1)$ additional steps for computing $\widehat{f}$. Thus, MUSCLE has a computation complexity given in (28). The term $\max_{0\leq k\leq\widehat{K}}\left(r_{k+1}-\min\mathcal{A}_{k}\right)^{3}$ in the right hand side of (28) depends on the underlying signal as well as the noise. For instance, if the signal consists of segments of bounded lengths and the noise is mild, then it happens most likely that $\widehat{K}=O(n)$ and also $\max_{0\leq k\leq\widehat{K}}\left(r_{k+1}-\min\mathcal{A}_{k}\right)^{3}=O(1)$. In this scenario, MUSCLE attains its best computational complexity of $O(n\log n)$. In general, as $r_{k}\leq n$, it holds always that $\max_{0\leq k\leq\widehat{K}}\left(r_{k+1}-\min\mathcal{A}_{k}\right)^{3}\leq n^{3}$. By (28) we obtain the asserted upper bound $O(n^{4}\log n)$ on the computational complexity of MUSCLE. The overall memory complexity is dominated by the storage of the wavelet tree, which is thus $O(n\log n)$.
@@ C.2 Proof of Theorem 6
Table 3: Comparison of different solutions to range quantile problem: By Naive, DH, MST, PST, and WT, we denote the solutions provided by naive implementation, double-heaps, merge sort tree, persistent segment tree, and wavelet tree, respectively. The table includes their space requirements, computational complexities of the preprocessing step, each query, and the total complexity for $m$ queries, where $m\leq n^{3}$. All results are concluded from (Castro et al., , 2016).
Methods Space Preprocessing Query Total

### v3

@@ Appendix C Implementation Details
The pseudocode of the proposed MUSCLE algorithm is provided in Algorithm 1.
@@
[…]
@@
15:         for $l=j-i,\dots,1$ do $\triangleright$ For dyadic intervals, $l\in\{2^{p},2^{p-1},\dots,1\}$, $p=\lfloor{{\log_{2}(j-i)}}\rfloor$
@@
32:           $S\leftarrow 1$ $\triangleright$ Pruning 1: drop all indices $j^{\prime}>j$, as no $j^{\prime}>j$ with $\mathcal{I}([i/n,j^{\prime}/n])=1$
@@
35:           break $\triangleright$ Pruning 2: drop all indices $i^{\prime}<i$, as no $i^{\prime}<i$ with $\mathcal{I}([i^{\prime}/n,j/n])=1$
@@ C.1 Computation of local tests
$T_{I}(Z,\theta)=\underset{J\subseteq I}{\max}\sqrt{2L_{J}(Z,\theta)}-\sqrt{2\log\frac{e\lvert I\rvert}{\lvert J\rvert}}$,
[…]
$\overline{W}_{J}\log\frac{\overline{W}_{J}}{\beta}+(1-\overline{W}_{J})\log\frac{1-\overline{W}_{J}}{1-\beta}\leq\frac{1}{2\lvert J\rvert}\,\left(q_{\alpha}(\lvert I\rvert)+\sqrt{2\log\frac{e\lvert I\rvert}{\lvert J\rvert}}\right)^{2}\;\eqqcolon\;\tilde{q}_{\lvert I\rvert,\lvert J\rvert}$. (25)
Here $\tilde{q}_{\lvert I\rvert,\lvert J\rvert}=\tilde{q}_{k,l}$ depends only on $k=\lvert I\rvert$ and $l=\lvert J\rvert$. The function $g_{\beta}:(0,1)\to\mathbb{R}$, $x\mapsto x\log\frac{x}{\beta}+(1-x)\log\frac{1-x}{1-\beta}$ is strictly convex on $(0,1)$ and has a unique minimizer at $x=\beta$. Let $u_{k,l}$ and $v_{k,l}$ with $u_{k,l}\leq v_{k,l}$ be the solutions of $g_{\beta}(x)=\tilde{q}_{k,l}$. Then the inequality in (25) becomes $\overline{W}_{J}\in[u_{k,l},v_{k,l}],\text{ for all intervals }J\subseteq I,\,k=\lvert I\rvert,l=\lvert J\rvert$. By the definition of $W_{i}$’s, the above displayed relation holds if and only if $\theta\in\bigcap_{J\subseteq I}\left[Z_{J,u_{k,l}},Z_{J,v_{k,l}}\right]$ where $Z_{J,q}$ is the $q$-quantile of $Z_{J}=(Z_{i})_{i\in J}$ for any $q\in(0,1)$. This intersection is empty if and only if there is no $\theta\in\mathbb{R}$ fulfils the multiscale constraints on $I$, in which case at least one additional change point is required. The subinterval $J$ is varying in $I$, and the corresponding quantities $u_{k,l}$ and $v_{k,l}$ are changing as well. It amounts to computing the so-called range quantiles, which is formulated as the range quantile problem below.
@@ 2. Output:
In computer science, the range quantile problem is solved in two steps: a preprocessing step and a query step. The preprocessing step transforms the original input $Y$ into a certain data structure, and in the query step each query is answered efficiently based on the data structure created in the preprocessing step. We focus on the data structure of wavelet tree and refer to Remark 5 and Table 4 for alternative data structures.
@@
Table 4: Comparison of different solutions to range quantile problem: By Naive, DH, MST, PST, and WT, we denote the solutions provided by naive implementation, double-heaps, merge sort tree, persistent segment tree, and wavelet tree, respectively. The table includes their space requirements, computational complexities of the preprocessing step, each query, and the total complexity for $m$ queries, where $m\leq n^{3}$. All results are from (Castro et al., , 2016).
Method Space Preprocessing Query Total
@@
A wavelet tree is a binary tree whose nodes consist of a value vector and a rank vector. The value vector of root is exactly the whole input and it is spitted into two children by comparing with the median of the value vector. More precisely, its left child contains all elements smaller than median while its right child consists of all elements larger or equal than median (for simplicity we assume that all $n$ data are distinct). For $i\in\{1,\dots,n\}$, the $i$-th entry of rank vector in root indicates the number of elements belonging to the left child between the first and $i$-th value. A wavelet tree can be constructed recursively until its leaves (i.e., nodes with only one element). Note that it is possible to compute all splitting required medians in $O(n\log n)$ steps and each layer can be constructed in $O(n)$ time. Thus, a wavelet tree is created in $O(n\log n)$ steps. Each single query $Q=(L,R,q)$ is answered in $O(\log n)$ steps as follows: Let $k=\lceil{(R-L+1)\cdot q}\rceil$ denote order of the $q$-th quantile of $(Y_{L},\dots,Y_{R})$. At each node we compute the difference between the rank of $R$ and rank of L1L-1, which equals to the number of elements in $(Y_{L},\dots,Y_{R})$ belonging to its left child. If it is greater or equal than $k$, then the target value is in the left child; otherwise, the target value is in the right child. We continue this process until the last layer and return value in the last leaf. The computation cost of at each node is $O(1)$, since all ranks are precomputed. Note the depth of a wavelet tree is $O(\log n)$, so the computational complexity of each query is $O(\log n)$.
This leads to the following proposition.
Proposition 12.
C.2 Proof of Theorem 6
We proceed in the following steps.
Recursion. Consider the recursive relation in (16). The subproblem with $j=1$ is solved immediately (cf. line 1 in Algorithm 1). For any $j\geq 2$, the $j$-th subproblem can be solved recursively by (16) via checking whether there is a constant function fulfilling the multiscale on $[i/n,j/n)$, i.e., $\mathcal{I}([i/n,j/n))=1$ (lines 12–27 in Algorithm 1). One should keep in mind that even if we knew $\mathcal{I}\left(\left[i/n,j/n\right)\right)=0$ for some $i$, we can not conclude $\mathcal{I}\left(\left[i^{\prime}/n,j/n\right)\right)=0$ for $i^{\prime}<i$, due to the varying penalties and quantiles. Since $\widehat{K}[j]$ requires the minimization among all indices $i\leq j-1$ with $\mathcal{I}([i/n,j/n))=1$, all indices should be taken into account.
Pruning. Nevertheless, the checking of some indices can be avoided if we employ some pruning techniques similarly as in Li et al., (2016). Towards this, we consider the test statistic $T_{I}^{0}(Z,\theta)={\max}_{J\subseteq I}\sqrt{2L_{J}(W,\theta)}-\sqrt{2\log\left({en}/{\lvert J\rvert}\right)}$, where the largest penalty with size $n$ is subtracted. Then it holds for all $Z$, $\theta$ and $I$ that $T_{I}(Z,\theta)\geq T_{I}^{0}(Z,\theta)$. During the minimization over indices ranging from i=j1i=j-1 to $i=0$, if we find an index $i$ with $\mathcal{I}\left(\left[i/n,j/n\right)\right)=0$, then we check whether there exists a $\theta\in\mathbb{R}$ such that $T_{[i/n,j/n]}^{0}(Z,\theta)\leq\max_{m\leq n}q_{\alpha}(m),$ which is the most relaxed multiscale constraint. If no constant $\theta$ fulfills this constraint displayed above, then $\mathcal{I}(I)=0$ for all $I$ containing $[i/n,j/n)$ and thus the indices $1,\dots,i-1$ can be excluded from further consideration (cf. lines 29–33 in Algorithm 1).
Final dynamic program. The final procedure is as follows. Let $\mathcal{A}_{0}=\{0\}$ and $\mathcal{B}_{0}=\{1,\dots,n\}$ and for $k=1\dots n$,
$r_{k}$ $\coloneqq\left\{j:\exists i\in\mathcal{A}_{k-1},\theta\in\mathbb{R}\text{ s.t. }T_{[\frac{i}{n},\frac{j}{n}]}^{0}(Z,\theta)\leq\underset{m}{\max}q_{\alpha}(m)\right\},$
$\mathcal{A}_{k}$ $\coloneqq\{j\in\mathcal{B}_{k-1}\cap[1,r_{k}]:\,\mathcal{I}([\tfrac{i}{n},\tfrac{j}{n}))=1\text{ for some }i\in\mathcal{A}_{k-1}\},$
$\mathcal{B}_{k}$ $\coloneqq\mathcal{B}_{k-1}\setminus\mathcal{A}_{k}.$
Complexity analysis. We first preprocess data $(Z_{1},\dots,Z_{n})$ by constructing a wavelet tree, which requires $O(n\log n)$ steps and $O(n\log n)$ memory (line 1 in Algorithm 1). Then, by Proposition 12, $\mathcal{I}([i/n,j/n))$ can be computed in $O((j-i)^{2}\log n)$ steps (lines 13–18 in Algorithm 1). Thus, the overall complexity for computing $\widehat{K}$ is, up to a constant factor,
$\sum_{k=0}^{\widehat{K}}\sum_{i\in\mathcal{A}_{k}}\sum_{j=i+1}^{r_{k+1}}(j-i)^{2}\log n\lesssim\log n\sum_{k=0}^{\widehat{K}}\lvert\mathcal{A}_{k}\rvert\left(r_{k+1}-\min\mathcal{A}_{k}\right)^{3}\lesssim n\log n\cdot\underset{0\leq k\leq\widehat{K}}{\max}\left(r_{k+1}-\min\mathcal{A}_{k}\right)^{3}.$ (26)
The MUSCLE $\widehat{f}$ can be computed long with $\widehat{K}$ in (7). When checking each index in the dynamic program, we need only $O(1)$ additional steps for $\widehat{f}$. Thus, MUSCLE has a computation complexity as in (26). The term $\max_{0\leq k\leq\widehat{K}}\left(r_{k+1}-\min\mathcal{A}_{k}\right)^{3}$ in (26) depends on the underlying signal and the noise. For instance, if the signal consists of segments of bounded lengths and the noise is mild, then it happens most likely that $\widehat{K}=O(n)$ and also $\max_{0\leq k\leq\widehat{K}}\left(r_{k+1}-\min\mathcal{A}_{k}\right)^{3}=O(1)$. In this scenario, MUSCLE attains its best computational complexity of $O(n\log n)$. In general, as $r_{k}\leq n$, it holds always that $\max_{0\leq k\leq\widehat{K}}\left(r_{k+1}-\min\mathcal{A}_{k}\right)^{3}\leq n^{3}$. By (26) we obtain the asserted upper bound $O(n^{4}\log n)$ on the computational complexity, which is reduced to $O(n^{3}\log n)$ if only intervals of dyadic lengths are considered. The overall memory complexity is dominated by the storage of the wavelet tree, which is $O(n\log n)$.